a)    Identify the iterated integrals which are of the following holding Fubini’s Theorem. Give mathematical reason.

i)   ∫_0^ln3▒∫_0^ln2▒〖e^(x+2y) dydx〗                                    iii)  ∫_3^5▒∫_1^2▒〖1/〖(x+y)〗^2 dydx〗

ii)       ∫_(-1)^1▒∫_(〖-x〗^2)^x▒〖(x^2-y)dydx〗                            iv)  ∫_1^(3⁄2)▒∫_y^(3-y)▒ydxdy


a)    The volume in the first octant bounded by the coordinate planes, the plane y = 4, and the

plane  (x/3) + (z/5) = 1.


b)    Sketch the solid in the first octant that is enclosed by the planes x = 0, z = 0, x = 5,

 z y = 0, and            z = −2y + 6.


1
Expert's answer
2021-02-01T10:34:47-0500

Solution

a)

NOTE: Identify the iterated integrals which are of the following holding Fubini’s Theorem. Give mathematical reason.


Fubini's Theorem holds only if a function is continous on a Rectangle "R=[a,b] \times [c,d]"


(i)


"\intop_0^{ln3} \intop_0^{ln2}e^{x+2y}dydx"

exponential function will be always continous. Also we have a rectangle


"[0, ln2]\times[0, ln3]"

(ii)


"\intop_{-1}^1\intop_{-x^2}^x (x^2-y)dydx"

The given case does not hold Fubini's Theorem since limits of y contain variable parts also. It should be continuos on the rectangle area.


(iii)


"\intop_{3}^5\intop_{1}^2\frac{1}{(x+y)^2}dydx"


The given function will be discontinous if the denominator equals to zero


"(x+y^2)=0\\
x+y=0\\
x=-y"

but for the given limits, x and y will not be negative. Thus x=-y is not possible and so it is continous on Rectangle.


(iv) Similar to the previous case, limits of x contain variable part, thus it does not hold Fubini's Theorem.


Hence

"\implies" (i), (iii) holds Fubini's Theorem, (ii), (iv) does not hold Fubini's Theorem


a)    The volume in the first octant bounded by the coordinate planes, the plane y = 4, and the

plane  (x/3) + (z/5) = 1.


"\frac z5=(1-\frac x3) \implies z=(\frac{15-5x}{3})\\

\therefore V= \iiint_{E}\delta v = \intop_{y=0}^4 \intop_{x=0}^3 \intop_{0}^{\frac{15-5x}{3}}\delta z \delta x \delta y\\

\therefore V= \intop_{0}^4 \intop_{0}^3 (\frac{15-5x}{3})\ \delta x \delta y = \intop_{0}^4 (\intop_{0}^3 \frac{15-5x}{3}|\ \delta x \delta y\\
= \intop_{0}^4 5x -\frac{5}{3}(\frac{x^2}{2})|_0^3\ \delta y = \intop_{0}^4 (15 -\frac{45}{6})\ \delta y\\
= (15 -\frac{45}{6})(4-0)=\frac{15}{2}(4)= 30\ units^3"


b)    Sketch the solid in the first octant that is enclosed by the planes x = 0, z = 0, x = 5,

 z y = 0, and           z = −2y + 6.